Question 41

The price of widgets has decreased from $6 to $5, causing the quantity demanded to increase from 200 to 220 units. Use the midpoint method to calculate the price elasticity of demand. Round to two decimal places and use absolute values.
Answer #1

Price elasticity of demand using mid point formula ÷
Ed= (Q2-Q1)/[(Q2+Q1)/2] ÷ (P2-P1)/[(P2+P1)/2]

Ed= (220-200)/[(220+200)/2] ÷ (5-6)/[(5+6)/2]

Ed= 20/[(420)/2] ÷ 1/[(11)/2]

We didn't use negative sign because absolute value is asked

Ed= 20/210 ÷ 1/5.5

Ed= 0.095÷0.1818

Ed= 0.5238

Ed= 0.52